# TGDataset: Collecting and Exploring the Largest Telegram Channels Dataset

### v2

@@ 6. Use cases
Violence and extremism. According to several newspapers, the Capitol Hill riot of January 2021 was planned months before also leveraging Telegram channels and groups (McEvoy, 2021; Heilweil, 2021). In response to this event, the platform has intensified its monitoring and obscured dozens of public channels promoting calls to violence (Hatmaker, 2021). Nonetheless, despite the commitment of Telegram to obfuscate these channels, within our dataset, there are several public channels promoting the spread of neo-Nazi ideologies or calls to violence (e.g., White Aryan Woman, Feuerkrieg Division **OFFICIAL**). Therefore, the problem is still far from being solved. The TGDataset can help study and characterize these channels. Indeed, the main problem with this type of channel is the absence of a dataset containing them, as they are obscured or blocked once discovered. Instead, the TGDataset includes several unblurred channels of this typology and, therefore, can be used to analyze their features and build a machine-learning model able to detect them automatically.
@@ 7. Ethical Considerations
This paper presents the TGDataset, a new dataset that includes 120,979 Telegram channels and over 400 million messages. The data collection was conducted with care to only include information from Telegram channels and exclude personal data such as usernames, phone numbers, and subscribed channels.
Our data collection process complies with Telegram’s Terms of Service (Telegram, 2024b) and Telegram API’s Terms of Service (Telegram, 2024a), as there are no clauses that prohibit the collection of public chat data. Additionally, we have taken care to avoid flooding the platform with requests during data collection.
Moreover, our research complies with the academic research exemptions outlined in Article 85 of the GDPR (TEXT, 2024), which provide certain flexibilities for processing publicly available data in the interest of freedom of expression and academic purposes. As our work exclusively involves publicly accessible content and does not process private user data, it falls under the legitimate or public interest exemptions. Furthermore, we adhere to the principle of data minimization, as emphasized in the GDPR’s guidelines (Europe, 2024): We limit our analysis to admin-generated public posts, ensuring that only the information strictly necessary for our research objectives is processed.
@@ 9. Conclusions and Future works
In this work, we present the TGDataset (Anonymous, 2024b), a collection of more than 120,000 public Telegram channels that, to the best of our knowledge, is the largest collection of channels publicly available. After characterizing the main quantitative aspect of the TGDataset, we performed language detection to understand which are the most popular languages on the dataset. Then, we investigate the main topics covered by the English channels. Additionally, we released as open source the script we used to analyze it and the labeling we obtained (language and topic of the channels). In this paper, we investigate a few possible use cases in which our dataset can be extremely useful. In particular, our preliminary study of the TGDataset revealed the presence on Telegram of channels spreading questionable content and conspiracy theories. Moreover, we observed the presence of several borderline activities and channels dealing with dubious ethical content. With the release of the TGDataset, we aim to provide a valuable resource to researchers that enables further investigation into these areas, leading to a more refined understanding of Telegram and helping to mitigate potential risks to users.

### v3

@@ 6. Use cases
Violence and extremism. According to several newspapers, the Capitol Hill riot of January 2021 was planned months before also leveraging Telegram channels and groups (McEvoy, 2021; Heilweil, 2021). In response to this event, the platform has intensified its monitoring and obscured dozens of public channels promoting calls to violence (Hatmaker, 2021). Nonetheless, despite the commitment of Telegram to obfuscate these channels, within our dataset, there are several public channels promoting the spread of neo-Nazi ideologies or calls to violence (e.g., White Aryan Woman, Feuerkrieg Division **OFFICIAL**). Therefore, the problem is still far from being solved. The TGDataset can help study and characterize these channels. Indeed, the main problem with these channels is the absence of a dataset containing them, as they are obscured or blocked once discovered. Instead, the TGDataset includes several unblurred channels of this typology that can be used to analyze their features and build a machine-learning model able to detect them automatically. This study requires the Message and Description fields to analyze harmful content and obtain context on the purposes of these channels, and the Forwarded from field to trace the channel behind it.
Temporal analyses and information propagation. The TGDataset represents a valuable resource for exploring the temporal dynamics of information spread by leveraging message timestamps and channels’ creation dates. Message timestamps (Date field in Text messages and Media messages) facilitate granular analyses of dissemination patterns, enabling researchers to track how content spreads within and across channels. Instead, channels’ creation dates (Creation date field) enable the analysis of the correlation between a channel’s longevity and its influence on information propagation. These temporal attributes support a variety of studies. For example, if combined with the forwarding information (Is forwarded and Forwarded from fields in Text messages and Media messages), they allow for tracing the rise and evolution of trending topics, assessing the velocity of information diffusion for different content types, identifying peak activity periods for specific topics or channels, and examining shifts in information flow patterns over time. Moreover, beyond individual channels, the communities identified in the TGDataset present opportunities for studying inter-community information diffusion. Investigating these larger network structures could provide valuable insights into how information propagates across interconnected groups within Telegram, unveiling the mechanisms that drive broader dissemination trends.
@@ 7. Ethical Considerations
This paper presents the TGDataset, a new dataset that includes 120,979 Telegram channels and over 400 million messages. The data collection was conducted carefully to only include information from Telegram channels and exclude personal data such as usernames, phone numbers, and subscribed channels.
Our data collection process complies with Telegram’s Terms of Service (Telegram, 2024b) and Telegram API’s Terms of Service (Telegram, 2024a), as there are no clauses that prohibit the collection of public chat data. Additionally, we have taken care to avoid flooding the platform with requests during data collection. Moreover, our research complies with the academic research exemptions outlined in Article 85 of the GDPR (TEXT, 2024), which provide certain flexibilities for processing publicly available data in the interest of freedom of expression and academic purposes. As our work exclusively involves publicly accessible content and does not process private user data, it falls under the legitimate or public interest exemptions. Furthermore, we adhere to the principle of data minimization, as emphasized in the GDPR’s guidelines (Europe, 2024): We limit our analysis to admin-generated public posts, ensuring that only the information strictly necessary for our research objectives is processed.
@@ 9. Conclusions and Future works
In this work, we present the TGDataset (La Morgia et al., 2024b), a collection of more than 120,000 public Telegram channels that, to the best of our knowledge, is the largest collection of channels publicly available. After characterizing the main quantitative aspect of the TGDataset, we performed language detection to understand which are the most popular languages on the dataset. Then, we investigate the main topics covered by the English channels. Additionally, we publicly released the script we used to analyze it and the labeling we obtained (language and topic of the channels). In this paper, we investigate a few possible use cases in which our dataset can be extremely useful. In particular, our preliminary study of the TGDataset revealed some Telegram channels spread questionable content and conspiracy theories. Moreover, we observed the presence of several borderline activities and channels dealing with dubious ethical content. With the release of the TGDataset, we aim to provide a valuable resource to researchers that enables further investigation into these areas, leading to a more refined understanding of Telegram and helping to mitigate potential risks to users.
@@
10. Acknowledgments
This work has been partially funded by projects: MUR National Recovery and Resilience Plan, SERICS (PE00000014); and ST3P (B83C24003210001) under the ”Young Researchers 2024-SoE” Program funded by the Italian Ministry of University and Research (MUR).